OpenStudy (anonymous):

Using the properties of exponents and radicals, design at least three different equivalent forms of x². You must show how each one can be simplified back to x² in two or more steps. Stretch your mind and get creative! Keep in mind that something too simple, like x • x would not be acceptable since it takes only one step to convert it to x².

OpenStudy (anonymous):

@Akeller27 and @callie2240

OpenStudy (anonymous):

One form is the square root of x to the fourth power: (x^4)^1/2 Another form is the fourth root of x to the eighth power: (x^8)^1/4 Another form is the eighth root of x to the sixteenth power: (x^16)^1/8 Each one can be simplified to x^2 by multiplying the exponent within the parenthesis with the external exponent.

OpenStudy (anonymous):

Alright I need to come up with some that need more then 1 step

OpenStudy (anonymous):

thanks for the examples though

OpenStudy (anonymous):

hmm how would I add another step onto it lol?

OpenStudy (anonymous):

Since you need radicals... \[(x ^{\frac{ 3 }{ 2 }}) (\frac{ 1 }{ \sqrt{x}}) (\sqrt{x}) (\sqrt{x})\]

OpenStudy (anonymous):

lost connection